f(x) = 23 – 2:02 + 15.rShow how to find the absolute max and min using this equation.
Answer #1

not exist Given RO)=313- 2x? +15% f(x) = x) 3_202 +15%) dx (63)- (2x2) (154) XP ds 30 ?_4X+15 2 -4x+15=o 3X 2 0 32 - Aise ts